OpenStudy (anonymous):

What is the sum of a 17–term arithmetic sequence where the first term is 6 and the last term is –90? Choices –813 –790 –765 –714

OpenStudy (mertsj):

\[\S=\frac{n}{2}(a _{1}+a _{n})\]

OpenStudy (anonymous):

-714 is the answer

OpenStudy (anonymous):

am i right ?

OpenStudy (mertsj):

\[\frac{17}{2}(6-90)=-714\]

OpenStudy (anonymous):

thats what i got so thank you i understand

hero (hero):

:D